# Quantum symmetries of noncommutative tori

### v1

@@ 1. Introduction
Fusion categories are algebraic objects that characterize generalized finite symmetries in many areas of quantum physics and mathematics. They are the structures underlying the standard invariants of finite depth subfactors, and characterize the algebraic structure of superselection sectors in quantum field theory. In both of these contexts, a fusion category “acts by bimodules” on a von Neumann (or C*) algebra $A$, generalizing the notion of a (cocycle) group action. This provides a framework for non-invertible quantum symmetries. More precisely, an action of a unitary fusion $\mathcal{C}$ on a C*-algebra is a C*-tensor functor from $\mathcal{C}$ to the C*-tensor category $\text{Bim}(A)$ of bimodules.111by bimodule of C*-algebra we mean a correspondence and we also assume our C*-algebras are unital
This perspective on quantum symmetries originated in Vaughan Jones’ theory of subfactors [Jon83]. Popa’s fundamental results on subfactors translate to a complete classification of fusion category actions on amenable factors [Pop90, Pop95]. Much less is known for fusion category actions on simple amenable C*-algebras, and the situation is expected to be significantly more complicated. Recently, however, there has been significant interest in this direction [Izu02, CHPJ24, AKK23, GPN23, EGP23, Izu24, Pac23] (see the introduction to [GS23] for a guide to recent progress on the closely related problem of classifying group actions on amenable C*-algebras). Using the machinery of subfactor theory, it is fairly straightforward to construct (and even classify in some instances) a wide variety of actions on AF-algebras [Ocn88, EK98, Pop90, CHPJ24] and Cuntz-Kreiger algebras [Izu98, Izu01]. However, there appears to be a lack of examples in the literature beyond these cases.
A natural family of stably finite amenable C*-algebras to consider beyond the AF case are noncommutative tori. Given a skew-symmetric real $n\times n$ matrix $\Theta$, $A_{\Theta}$ is the universal C*-algebra generated by $n$ unitaries $U_{1},\dots U_{n}$ satisfying $U_{k}U_{j}=e^{2\pi i\Theta_{j,k}}U_{j}U_{k}$. This is the algebra of continuous functions on the noncommutative n-torus. If the $\mathbbm{R}^{\times}$ valued bicharacter $b_{\Theta}(x,y):=e^{2\pi i\langle x,\Theta y\rangle}$ on $\mathbbm{Z}^{n}$ is nondegenerate, then $A_{\Theta}$ is a simple nuclear AT C*-algebra with unique trace [Phi06]. The algebras $A_{\Theta}$ (and their smooth subalgebras) have wide-ranging applications in both condensed matter [BvESB94, Bel86] and high-energy physics [CDS98, DN01], and are the paradigmatic example in Connes’ noncommutative geometry [Con94].
@@
The fusion categories $SU(2)_{k}$ with $k+2$ an odd, square-free integer, and $k+1$ not a power of $2$ admit no action on any noncommutative torus of any rank.
@@
The basic idea we exploit in this paper comes from the results of [EE93, Phi06], which demonstrate that simple non-commutative tori are inductive limits of C*-algebras of the form $C(\mathbbm{T},A)$, where $A$ is a finite dimensional C*-algebra. In other words, simple noncommutative tori are AT-algebras. Our strategy consists of building actions of fusion categories on “building block” C*-algebras $C(\mathbbm{T},A)$, and then choosing sufficiently nice connecting maps between these blocks to build an inductive limit action on an AT-algebra, which can be identified with a noncommutative torus via general classification results. In general, we call actions built from circle algebras in this way AT-actions 2.9. The technical result we will use to produce our actions is summarized in Theorem 2.13, which gives a construction of an AT action from some basic categorical data.
@@ 2. Actions of unitary tensor categories on C*-algebras
Associated to a unital C*-algebra $A$ is a C*-tensor category $\text{Bim}(A)$. It consists of all (right) A-A C*-correspondences, which are finitely generated projective as right Hilbert modules. The relative tensor product is denoted $\boxtimes_{A}$. See [CHPJP22, Section 2.2] and references therein for definitions and discussion related to C* correspondences.
@@ 2.1. Inductive limit actions
One way to build an action of a fusion category ${\mathcal{C}}$ on a C*-algebra $A$ is to realize $A$ as the inductive limit of a sequence of C*-algebras $A_{1}\rightarrow_{\iota_{1}}A_{2}\rightarrow_{\iota_{2}}\cdots\rightarrow A$, and build actions at each level. We then equip the connecting maps with “${\mathcal{C}}$-equivariant” structures. This will yield an action of ${\mathcal{C}}$ on $A$ as in [CHPJ24, Proposition 4.4], [GPN23, Section 4].
@@
Note that $B^{{\mathcal{C}}}\subseteq B$ is a unital subalgebra and that $[B^{{\mathcal{C}}},\iota(A)]=0$. Indeed, we will think of this as a generalized “relative commutant” of the whole included action $j(\mathcal{C})$.
@@
This is positive definite on the algebraic tensor product, and thus we complete $\widetilde{F}(X)$ in the induced norm. In this paper, either $F(X)$ or $A$ will be finite-dimensional and thus no completion is required. Furthermore, for $f\in\mathcal{C}(X,Y)$, we can define
@@
    $\iota\otimes\text{id}_{C}=\nu|_{A\otimes 1_{C}}$
@@
We now recall some definitions from [CHPJ24] (c.f. [GPN23]).
@@
Let ${\mathcal{C}}$ be a fusion category and $A_{1}\rightarrow_{\iota_{1}}A_{2}\rightarrow_{\iota_{2}}\cdots$ be a tower of unital inclusions of C*-algebras. Inductive limit action data consists of a family of actions ${\mathcal{C}}\overset{F_{n}}{\curvearrowright}A_{n}$, and for each $n$, a $\mathcal{C}$-equivariant structure $j{n}$ on the homomorphism $\iota_{n}:A_{n}\rightarrow A_{n+1}$.
@@ 2.2. Extending simple stationary actions
Given a unitary fusion category $\mathcal{C}$ and a right finitely semisimple $\mathcal{C}$-module category $\mathcal{M}$ (for definitions see [REFs]), recall that the dual category $\mathcal{C}^{*}_{\mathcal{M}}=\text{End}_{\mathcal{C}}(\mathcal{M})^{mp}$ is the monoidal opposite of the (multi)-fusion category of ${\mathcal{C}}$-module endofunctors on $\mathcal{M}$. For example, the “trivial” ${\mathcal{C}}$-module category (${\mathcal{C}}$-acting on itself by right tensoring), the dual category is $\mathcal{C}$ itself.
@@
Stationary AF-actions. Given a unitary fusion category $\mathcal{C}$ with tensor product denoted $\boxtimes$, a right finitely semisimple $\mathcal{C}$-module category $\mathcal{M}$, an object $X\in\mathcal{C}^{*}_{\mathcal{M}}$ and a choice of $m\in\mathcal{M}$ containing every isomorphism class of simple in $\mathcal{M}$ as a summand, we can define an AF-action of $\mathcal{C}$ as follows (see also [EPJ24, Section 5]):
Set $L_{n}:=X^{\boxtimes n}\boxtimes m\in\mathcal{M}$. Define the finite dimensional C*-algebras $A_{n}:=\text{End}_{\mathcal{M}}(L_{n})$. Then we have an action $F_{n}:\mathcal{H}\rightarrow\text{Bim}(A_{n})$ given by
@@
The tensorators are easy to write down. Note this is simply the action of $\mathcal{C}$ on $\mathcal{M}\cong\text{Mod}(\text{End}(X^{n}\otimes n))$, translated to the bimodule formalism (as in Remark 2.2). The algebra connecting maps $\iota_{n}:A_{n}\rightarrow A_{n+1}$ are defined by $\iota_{n}(a):=1_{X}\otimes a$, and the bimodule connecting maps $j^{Y}_{n}:F_{n}(Y)\rightarrow F_{n+1}(Y)$ are defined by $j^{Y}_{n}(f):=1_{X}\otimes f$.
@@
The assumption that $X$ is a strong tensor generator guarantees that the AF algebra $A$ is simple and has a unique trace. The indecomposability condition is mostly for convenience, since it ensures $C^{*}(\mathcal{M})$ is unitary fusion rather than unitary multi-fusion. This assumption can almost certainly be removed with care, but we find it convenient to have a simple unit in $C^{*}_{\mathcal{M}}$ for the proofs below.
@@
$W(z):=z\otimes(e_{1}\circ e^{*}_{n})+\sum^{l}_{i=1}1_{C(\mathbbm{T})}\otimes(e_{i}\circ e^{*}_{i-1})\ \ +\sum_{x\in\text{Irr}1_{C(\mathbbm{T})}\otimes(\mathcal{D})-\{\mathbbm{1}\}}1_{C(\mathbbm{T})}\otimes Z_{x}$
@@
If the fusion matrix for $X$ on $\mathcal{M}$ is in $GL_{\text{rank}(\mathcal{M})}(\mathbbm{Z})$, then $K_{0}(A)\cong\mathbbm{Z}^{\text{rank}(\mathcal{M})}$ as abelian groups. Then one could take the AF-action of $\mathcal{C}$ on the AF-algebra $A$ built from the above data and simply tensor with the classifiable C*-algebra $D$ with $K_{0}\cong K_{1}\cong\mathbbm{Z}$ with unique trace. Then by the Kunneth product formula ([RS87]) and the classification theorem for simple nuclear C*-algebras [GLN14, EGLN15], the algebra $B\cong D\otimes A$. In particular, the tensor extension of the action on $A$ to $A\otimes D$ yields an action of $\mathcal{C}$ on $B$, which has the nice regularity property of being manifestly $Z$-stable in the sense of [EPJ24]. However, it was communicated to us by George Elliott that $D$ is not itself AT, so it is not clear whether the action constructed this way is an AT-action. This raises a question that could serve as a first goal for any attempt to classify fusion category actions on C*-algebras: are the actions on the AT-algebras constructed via the above theorem above equivalent to the ones obtained by tensoring the underlying AF-action by $D$?
@@ 3. Actions on noncommutative tori
The matrix $\Theta$ is said to be degenerate if there exists an $x\in\mathbbm{Q}^{n}$ such that $\langle x,\theta y\rangle\in\mathbbm{Q}$ for all $y\in\mathbbm{Q}^{n}$. Then by [Phi06], if $\Theta$ is non-degenerate than $A_{\Theta}$ is a simple AT C*-algebra with unique trace and real rank 00, which gives us access to Elliot’s classification theorem [Ell93]. This allows us to build C*-algebras isomorphic to $B$ by constructing a simple AT algebra with with the appropriate K-theory and traces.
@@
In this paper, the goal is to construct actions of interesting fusion categories on the topological noncommutative torus $A_{\Theta}$. We say an action of a tensor category $\mathcal{C}$ on $A_{\Theta}$ is algebraic if it restricts to an action on $A^{0}_{\Theta}$.
@@ 3.1. No-go results
Let $\mathcal{C}$ be a fusion category containing an object with non-integral dimension. Then there is no algebraic action of $\mathcal{C}$ on any $A_{\Theta}$ of any dimension.
@@
Suppose $\mathcal{C}$ is a unitary fusion whose dimensions are rationally independent. Then if $\text{rank}(\mathcal{C})$ is not a power of 2, there is no action of $\mathcal{C}$ on any noncommutative torus.
@@
By hypothesis, the dimension function $\text{FPdim}:\mathbbm{Q}[\mathcal{C}]\rightarrow\mathbbm{R}$ is a ring isomorphism onto its image. But since the set $\text{Irr}(\mathcal{C})$ forms a basis for the rational fusion ring $\mathbbm{Q}[\mathcal{C}]$ and the dimensions of these objects are algebraic integers [EGNO15, Proposition 3.3.4], $\mathbbm{Q}[\mathcal{C}]$ is a field. Therefore $K:=\mathbbm{Q}[\mathcal{C}]$ is a field, and $[K:\mathbbm{Q}]=\text{rank}(\mathcal{C})$. Then, if $\mathcal{C}$ had an action on the noncommutative torus $A_{\Theta}$ of rank $n$, this would induce a $\mathbbm{Z}[\mathcal{C}]$-module structure on $K_{0}(A_{\Theta})\cong\mathbbm{Z}^{2^{n-1}}$. Extending scalars gives a $\mathbbm{Q}[\mathcal{C}]$-module structure on $K_{0}(A_{\Theta})\otimes_{\mathbbm{Z}}\otimes\mathbbm{Q}\cong\mathbbm{Q}^{2^{n-1}}$. But as pointed out above, $\mathbbm{Q}[\mathcal{C}]$ is a field, so this module is free, i.e. as $\mathbbm{Q}[\mathcal{C}]$-modules,
@@
$Y_{a}\otimes Y_{j}\cong\oplus N^{c}_{ab}Y_{c}$
$N^{c}_{ab}=\delta_{0\leq c\leq\text{min}\{l-(a+b),a+b\}}$
@@
Suppose $k>1$ and $k+2$ is odd and square-free. Then the dimensions of simple objects of $\text{PSU}(2)_{k}$ are rationally independent.
@@
Now, since $q=e^{\frac{\pi i}{k+2}}$, $[m]_{q}=[k+2-m]_{q}$ for $m=0,1,\dots\frac{k+1}{2}$. This implies that the set of dimensions
@@
But for $m$ odd $q^{m}$ all of the roots of unity appearing in the above sum are distinct primitive roots of unity (see, for example [ht]), and since $k+2$ is odd and square-free, hence $2(k+2)$ is square-free hence the primitve $2(k+2)^{th}$ roots of unity are rationally independent. ∎
@@
Let $k>1$ such that $k+2$ is odd and square-free and $k+1\neq 2^{n}$. Then any unitary fusion category containing $\text{PSU}(2)_{k}$ as a full subcategory admits no topological action on a noncommutative torus of any rank. In particular, for these values of $k$, $\text{SU}(2)_{k}$ do not admit actions on noncommutative tori.
We will see in a later section examples of actions of $\text{PSU}(2)_{15}$ on noncommutative tori, which satisfy $15+1=2^{4}$.

### v3

@@ 1. Introduction
Fusion categories are algebraic objects that characterize generalized finite symmetries in many areas of quantum physics and mathematics. They are the structures underlying the standard invariants of finite depth subfactors, and characterize the algebraic structure of superselection sectors in quantum field theory. In both of these contexts, a fusion category “acts by bimodules” on a von Neumann (or C*) algebra $A$, generalizing the notion of a (cocycle) group action. This provides a framework for non-invertible quantum symmetries. More precisely, an action of a unitary fusion category $\mathcal{C}$ on a C*-algebra $A$ is a C*-tensor functor from $\mathcal{C}$ to the C*-tensor category $\text{Bim}(A)$ of bimodules.111by bimodule of C*-algebra we mean a correspondence and we also assume our C*-algebras are unital
This perspective on quantum symmetries originated in Vaughan Jones’ theory of subfactors [Jon83]. Popa’s fundamental results on subfactors translate to a complete classification of fusion category actions on amenable factors [Pop90, Pop95]. Much less is known for fusion category actions on simple amenable C*-algebras, and the situation is expected to be significantly more complicated. Recently, however, there has been significant interest in this direction [Izu02, CHPJ24, AKKOA, GPNOA, EGP23, IzuOA, PacOA] (see the introduction to [gabe2023dynamical] for a guide to recent progress on the closely related problem of classifying group actions on amenable C*-algebras). Using the machinery of subfactor theory, it is fairly straightforward to construct (and even classify in some instances) a wide variety of actions on AF-algebras [Ocn88, EK98, Pop90, CHPJ24] and Cuntz-Kreiger algebras [Izu98, Izu01]. However, there appears to be a lack of examples in the literature beyond these cases.
A natural family of stably finite amenable C*-algebras to consider beyond the AF case are noncommutative tori. Given a skew-symmetric real $n\times n$ matrix $\Theta$, $A_{\Theta}$ is the universal C*-algebra generated by $n$ unitaries $U_{1},\dots U_{n}$ satisfying $U_{k}U_{j}=e^{2\pi i\Theta_{j,k}}U_{j}U_{k}$. This is the algebra of continuous functions on the noncommutative n-torus. If the $\mathbbm{R}^{\times}$ valued bicharacter $b_{\Theta}(x,y):=e^{2\pi i\langle x,\Theta y\rangle}$ on $\mathbbm{Z}^{n}$ is nondegenerate, then $A_{\Theta}$ is a simple nuclear AT C*-algebra with unique trace [Phi83]. The algebras $A_{\Theta}$ (and their smooth subalgebras) have wide-ranging applications in both condensed matter [BvESB94, Bel86] and high-energy physics [CDS98, DN01], and are the paradigmatic example in Connes’ noncommutative geometry [Con94].
@@
The fusion categories $SU(2)_{k}$ with $k+2$ prime and $k+1$ not a power of $2$ admit no action on any noncommutative torus of any rank.
@@
The basic idea we exploit in this paper comes from the results of [EE93, Phi83], which demonstrate that simple non-commutative tori are inductive limits of C*-algebras of the form $C(\mathbbm{T},A)$, where $A$ is a finite dimensional C*-algebra. In other words, simple noncommutative tori are AT-algebras. Our strategy consists of building actions of fusion categories on “building block” C*-algebras $C(\mathbbm{T},A)$, and then choosing sufficiently nice connecting maps between these blocks to build an inductive limit action on an AT-algebra, which can be identified with a noncommutative torus via general $K$-theoretic classification results for AT-algebras. In particular, Elliott [Ell97, Ell95] gave a complete classification of simple AT-algebras through the ordered $K_{0}$-group, $K_{1}$, the simplex of tracial states and its pairing with $K_{0}$, which we can apply in our setting. In general, we call actions built from circle algebras in this way AT-actions 2.9. The technical result we will use to produce our actions is summarized in Theorem 2.13, which gives a construction of an AT action from some basic categorical data.
@@ 2. Actions of unitary tensor categories on C*-algebras
Associated to a unital C*-algebra $A$ is a C*-tensor category $\text{Bim}(A)$. It consists of all (right) $A$-$A$ C*-correspondences, which are finitely generated projective as right Hilbert modules. The relative tensor product is denoted $\boxtimes_{A}$. See [CHPJP22, Section 2.2] and references therein for definitions and discussion related to C*-correspondences.
@@
  • For a C-algebra $A$, an $A$-$A$ bimodule of finite Jones index type yields an element of the Kasparov group $KK(A,A)$ and so an element of $\text{End}(K_{0}(A))$ [KW00]. Hence an action of $\mathcal{C}$ on $A$, yields a unital homomorphism from $\text{Fus}(\mathcal{C})$ to $\text{End}(K_{0}(A))$.
@@ 2.1. Inductive limit actions
One way to build an action of a fusion category ${\mathcal{C}}$ on a C*-algebra $A$ is to realize $A$ as the inductive limit of a sequence of C*-algebras $A_{1}\rightarrow_{\iota_{1}}A_{2}\rightarrow_{\iota_{2}}\cdots\rightarrow A$, and build actions at each level. We then equip the connecting maps with “${\mathcal{C}}$-equivariant” structures. This will yield an action of ${\mathcal{C}}$ on $A$ as in [CHPJ24, Proposition 4.4], [GPNOA, Section 4].
@@
Note that $B^{{\mathcal{C}}}_{\iota}\subseteq B$ is a unital subalgebra and that $[B^{{\mathcal{C}}},\iota(A)]=0$. Indeed, we will think of this as a generalized “relative commutant” of the whole included action $j(\mathcal{C})$.
@@
This is positive definite on the algebraic tensor product, and thus we complete $\widetilde{F}(X)$ in the induced norm. In this paper, either $F(X)$ or $C$ will be finite-dimensional and thus no completion is required. Furthermore, for $f\in\mathcal{C}(X,Y)$, we can define
@@
    $\iota\otimes\text{id}_{C}=\nu|_{A\otimes{C}}$
@@
We now recall some definitions from [CHPJ24] (c.f. [GPNOA]).
@@
Let ${\mathcal{C}}$ be a fusion category and $A_{1}\rightarrow_{\iota_{1}}A_{2}\rightarrow_{\iota_{2}}\cdots$ be a tower of unital inclusions of C*-algebras. Inductive limit action data consists of a family of actions ${\mathcal{C}}\overset{F_{n}}{\curvearrowright}A_{n}$, and for each $n$, a $\mathcal{C}$-equivariant structure $j_{n}$ on the homomorphism $\iota_{n}:A_{n}\rightarrow A_{n+1}$.
@@ 2.2. Extending simple stationary actions
Given a unitary fusion category $\mathcal{C}$ and a right finitely semisimple $\mathcal{C}$-module category $\mathcal{M}$ (for definitions see [JP17]), recall that the dual category $\mathcal{C}^{*}_{\mathcal{M}}=\text{End}_{\mathcal{C}}(\mathcal{M})^{mp}$ is the monoidal opposite of the (multi)-fusion category of ${\mathcal{C}}$-module endofunctors on $\mathcal{M}$. For example, the “trivial” ${\mathcal{C}}$-module category (${\mathcal{C}}$-acting on itself by right tensoring), the dual category is $\mathcal{C}$ itself.
@@
Stationary AF-actions. Given a unitary fusion category $\mathcal{C}$ with tensor product denoted $\boxtimes$, a right finitely semisimple $\mathcal{C}$-module category $\mathcal{M}$, an object $X\in\mathcal{C}^{*}_{\mathcal{M}}$ and a choice of $m\in\mathcal{M}$ containing every isomorphism class of simple in $\mathcal{M}$ as a summand, we can define an AF-action of $\mathcal{C}$ as follows (see also [EPJOA, Section 5]):
Set $L_{n}:=X^{\boxtimes n}\triangleright m\in\mathcal{M}$. Define the finite dimensional C*-algebras $A_{n}:=\text{End}_{\mathcal{M}}(L_{n})$. Then we have an action $F_{n}:\mathcal{C}\rightarrow\text{Bim}(A_{n})$ given by
@@
The tensorators are easy to write down. Note this is simply the action of $\mathcal{C}$ on $\mathcal{M}\cong\text{Mod}(\text{End}(X^{n}\boxtimes m))$, translated to the bimodule formalism (as in Remark 2.2). The algebra connecting maps $\iota_{n}:A_{n}\rightarrow A_{n+1}$ are defined by $\iota_{n}(a):=1_{X}\otimes a$, and the bimodule connecting maps $j^{Y}_{n}:F_{n}(Y)\rightarrow F_{n+1}(Y)$ are defined by $j^{Y}_{n}(f):=1_{X}\otimes f$.
@@
The assumption that $X$ is a strong tensor generator guarantees that the AF algebra $A$ is simple and has a unique trace. The indecomposability condition is mostly for convenience, since it ensures $C^{*}_{\mathcal{M}}$ is unitary fusion rather than unitary multi-fusion. This assumption can almost certainly be removed with care, but we find it convenient to have a simple unit in $C^{*}_{\mathcal{M}}$ for the proofs below.
@@
$W(z):=z\otimes(e_{1}\circ e^{*}_{n})+\sum^{l}_{i=1}1_{C(\mathbbm{T})}\otimes(e_{i}\circ e^{*}_{i-1})\ \ +\sum_{x\in\text{Irr}(\mathcal{D})-\{\mathbbm{1}\}}1_{C(\mathbbm{T})}\otimes Z_{x}$
@@
If the fusion matrix for $X$ on $\mathcal{M}$ is in $GL_{\text{rank}(\mathcal{M})}(\mathbbm{Z})$, then $K_{0}(A)\cong\mathbbm{Z}^{\text{rank}(\mathcal{M})}$ as abelian groups. Then one could take the AF-action of $\mathcal{C}$ on the AF-algebra $A$ built from the above data and simply tensor with the classifiable C*-algebra $D$ with $K_{0}\cong K_{1}\cong\mathbbm{Z}$ with unique trace. Then by the Kunneth product formula ([RS87]) and the classification theorem for simple nuclear C*-algebras [GLN14, EGLN15], the algebra $B\cong D\otimes A$. In particular, the tensor extension of the action on $A$ to $A\otimes D$ yields an action of $\mathcal{C}$ on $B$, which has the nice regularity property of being manifestly $Z$-stable in the sense of [EPJOA]. However, it was communicated to us by George Elliott that $D$ is not itself AT, so it is not clear whether the action constructed this way is an AT-action. This raises a question that could serve as a first goal for any attempt to classify fusion category actions on C*-algebras: are the actions on the AT-algebras constructed via the above theorem above equivalent to the ones obtained by tensoring the underlying AF-action by $D$?
@@ 3. Actions on noncommutative tori
The matrix $\Theta$ is said to be degenerate if there exists an $x\in\mathbbm{Q}^{n}$ such that $\langle x,\Theta y\rangle\in\mathbbm{Q}$ for all $y\in\mathbbm{Q}^{n}$. Then by [Phi83], if $\Theta$ is non-degenerate then $A_{\Theta}$ is a simple AT C*-algebra with unique trace and real rank 00, which gives us access to Elliot’s classification theorem [Ell93]. This allows us to build C*-algebras isomorphic to $B$ by constructing a simple AT algebra with the appropriate K-theory and traces.
@@
In this paper, the goal is to construct actions of interesting fusion categories on the topological noncommutative torus $A_{\Theta}$.
@@ 3.1. No-go results
Let $\mathcal{C}$ be a fusion category containing an object with non-integral dimension, and let $A^{0}_{\Theta}$ be an algebraic noncommutative torus of any dimenison. There is no linear monoidal functor $\mathcal{C}\rightarrow\text{Bim}(A^{0}_{\Theta})$222in this context, Bim denotes the linear monoidal category of purely algebraic bimodules with relative tensor product.
@@
Suppose $\mathcal{C}$ is a unitary fusion category whose dimensions of simple objects are rationally independent. Then if $\text{rank}(\mathcal{C})$ is not a power of 2, there is no action of $\mathcal{C}$ on any noncommutative torus.
@@
By hypothesis, the dimension function $\text{FPdim}:\mathbbm{Q}[\mathcal{C}]\rightarrow\mathbbm{R}$ is a ring isomorphism onto its image. But since the set $\text{Irr}(\mathcal{C})$ forms a basis for the rational fusion ring $\mathbbm{Q}[\mathcal{C}]$ and the dimensions of these objects are algebraic integers [EGNO15, Proposition 3.3.4], $\mathbbm{Q}[\mathcal{C}]$ is a field. Therefore $K:=\mathbbm{Q}[\mathcal{C}]$ is a field, and $[K:\mathbbm{Q}]=\text{rank}(\mathcal{C})$. Then, if $\mathcal{C}$ had an action on the noncommutative torus $A_{\Theta}$ of rank $n$, this would induce a $\mathbbm{Z}[\mathcal{C}]$-module structure on $K_{0}(A_{\Theta})\cong\mathbbm{Z}^{2^{n-1}}$. Extending scalars gives a $\mathbbm{Q}[\mathcal{C}]$-module structure on $K_{0}(A_{\Theta})\otimes_{\mathbbm{Z}}\mathbbm{Q}\cong\mathbbm{Q}^{2^{n-1}}$. But as pointed out above, $\mathbbm{Q}[\mathcal{C}]$ is a field, so this module is free, i.e. as $\mathbbm{Q}[\mathcal{C}]$-modules,
@@
$Y_{a}\otimes Y_{b}\cong\oplus N^{c}_{ab}Y_{c}$
$N^{c}_{ab}=\delta_{|a-b|\leq c\leq\text{min}\{l-(a+b),a+b\}}$
@@
Suppose $k>1$ and $k+2$ is prime. Then the dimensions of simple objects of $\text{PSU}(2)_{k}$ are rationally independent.
@@
Since $q=e^{\frac{\pi i}{k+2}}$, $[m]_{q}=[k+2-m]_{q}$ for $m=0,1,\dots\frac{k+1}{2}$. This implies that the set of dimensions
@@
The result now follows, for example see [Trile].
@@
Let $k>1$ such that $k+2$ is prime and $k+1\neq 2^{n}$. Then any unitary fusion category containing $\text{PSU}(2)_{k}$ as a full subcategory admits no topological action on a noncommutative torus of any rank. In particular, for these values of $k$, $\text{SU}(2)_{k}$ do not admit actions on noncommutative tori.
We will see in a later section examples of actions of $\text{PSU}(2)_{15}$ on noncommutative tori. Note $15+2$ is prime, but $15+1=2^{4}$.